OpenStudy (anonymous):

Simplify \[5^{-2}\] by making the exponent positive

OpenStudy (anonymous):

@mathstudent55 hey come

OpenStudy (anonymous):

there you are <3

OpenStudy (mathstudent55):

\( a^{-n} = \dfrac{1}{a^n} \)

OpenStudy (mathstudent55):

That is how you change a negative exponent to a positive exponent.

OpenStudy (anonymous):

so would it be 1 over 5^2?

OpenStudy (mathstudent55):

That's all there is to it.

OpenStudy (anonymous):

thanks i got more questions

OpenStudy (anonymous):

i got a really hard one hun i have no idea what it is and im stressing over it :(

OpenStudy (mathstudent55):

\( a^{-n} = \dfrac{1}{a^n}\) \(5^{-2} = \dfrac{1}{5^2} \)